Determine the standard form of the equatiion of the line that contains (-6,-6) and (-5,-2)
Answer by stanbon(48568)

Find the slope: (-6-(-2))/(-6-(-5))
=-4/-1=4
Use one of the poiints and the slope in the form y=mx+b to find "b".
Let's use the point (-6,-6)
y=-6,x=-6,m=4
-6=4(-6)+b
-6=-24+b
b=18
The equation in slope-intercept form is y=4x+18
Standard form is ax+bx=c
So, 4x-y=-18
Cheers,
Stan H.